OpenStudy (anonymous):

a motor car covers 1/3 part of total distance with velocity V1 = 10km/hr and second 1/3 part with velocity V2 = 20km/hr and rest 1/3 part with velocity V3 = 60km/hr . What is the average speed of the car ?

OpenStudy (anonymous):

Average speed of the car = (V1+V2+V3)/3

OpenStudy (anonymous):

now can u?

OpenStudy (anonymous):

Avg speed = 10+20+60 / 3 = 90/3 = 30km/hr correct ?

OpenStudy (anonymous):

very correct

OpenStudy (anonymous):

u got it?

OpenStudy (anonymous):

why the velocities / 3 ?

OpenStudy (anonymous):

Actually it was like this: Average speed of the car = 1/3 V1 + 1/3 V2 + 1/3 V3 = (V1+V2+V3)/3

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

oh okay .. thanks :)